The function F : R^2 → R defined by F(x,y)= (y + 2 , x + y) is locally invertible at any (x, y)∈R^2.
1
Expert's answer
2020-03-09T11:31:43-0400

The function is locally invertible if the Jacobian is not zero.

"\begin{vmatrix}
 (y+20)_x & (x+y)_x \\
 (y+2)_y & (x+y)_y
\end{vmatrix}
=
\begin{vmatrix}
 0 & 1 \\
 1 & 1
\end{vmatrix}
=0-1=-1\ne 0 \forall x,y\in R" Thus, the function is locally invertible.